An aspirin tablet contains 0.325g of acetylsalicylic acid (C9H8O4).

How many acetylsalicylic acid molecules does it contain?

What is the total number of atoms in that sample?
1
Expert's answer
2019-04-08T07:57:03-0400

(i) Find number of moleculs of "C_9H_8O_4"

"n = \frac{m}{M}"


"n(C_9H_8O_4) = \frac{0.325 g}{180.16 \frac{g}{mol}} = 0.00180 mol"


"N = N_A\times n"


"N(C_9H_8O_4) = 6.022\times 10^{23} mol^{-1} \times 0.00180 mol = 1.086\times10^{21} molecules"


(ii) find total number of atoms:


As 1 molecule of "C_9H_8O_4" contains 9+8+4 = 21 atoms, then


"N_{total}(atoms) = 21\times N(C_9H_8O_4) = 21\times 1.086\times10^{21} = 2.281\times 10^{22} atoms"